### v1

@@ 4. Half-Plane Property
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ be a univariate strictly-CLC polynomial for $t\geq 0$ with positive $a_{k}$ for $k=0,\dots,d$. For $1\leq d\leq 4$, then $f(t)$ is Hurwitz-stable polynomial.
@@
$\frac{a_{3}-a_{5}k_{1}}{a_{2}-a_{4}k_{1}}<\frac{a_{3}-a_{5}k_{2}}{a_{2}-a_{4}k_{2}}$ if and only if either $a_{2}a_{5}>a_{3}a_{4}$ if $k_{1}>k_{2}$, or $a_{2}a_{5}<a_{3}a_{4}$ if $k_{1}<k_{2}$.
@@
It follows from the following equivalence.
$\frac{a_{3}-a_{5}k_{1}}{a_{2}-a_{4}k_{1}}>\frac{a_{3}-a_{5}k_{2}}{a_{2}-a_{4}k_{2}}\Leftrightarrow a_{2}a_{5}(k_{1}-k_{2})<a_{3}a_{4}(k_{1}-k_{2})\Leftrightarrow\begin{cases}a_{2}a_{5}>a_{3}a_{4}\ \mbox{if}\ k_{1}>k_{2},\\ a_{2}a_{5}<a_{3}a_{4}\ \mbox{if}\ k_{1}<k_{2}\end{cases}$
Lemma 4.17.
For any positive $k$, $\frac{a_{1}}{a_{0}}>\frac{a_{3}-a_{5}k}{a_{2}-a_{4}k}$ if and only if $k<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}}$.
@@
This is obtained by applying the identity $\frac{a}{b}>\frac{c}{d}\Leftrightarrow ad>bc$. ∎
@@
Let $f(t)=\sum_{k=0}^{5}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,5$. If $a_{k}^{2}>\frac{k+1}{k}a_{k-1}a_{k+1}$ and $(a_{1}a_{4}-a_{0}a_{5})^{2}<(a_{1}a_{2}-a_{0}a_{3})(a_{3}a_{4}-a_{2}a_{5})$ for all $1\leq k\leq 4$, then $f(t)$ is Hurwitz-stable polynomial.
@@
By Theorem 4.2 and Remark 4.3, if $d=5$, it is sufficient to show that $\Delta_{k}>0$ for all $1\leq k\leq 4$ for ${\mathrm{Hur}}_{f(t)}=\begin{bmatrix}a_{1}&a_{3}&a_{5}&0&0\\ a_{0}&a_{2}&a_{4}&0&0\\ 0&a_{1}&a_{3}&a_{5}&0\\ 0&a_{0}&a_{2}&a_{4}&0\\ 0&0&a_{1}&a_{3}&a_{5}\end{bmatrix}$. By the given hypothesis, we have $a_{1}>0,a_{1}a_{2}-a_{0}a_{3}>0$ and $\begin{vmatrix}a_{1}&a_{3}&0\\ a_{0}&a_{2}&a_{4}\\ 0&a_{1}&a_{3}\end{vmatrix}>0$, which implies that $\begin{vmatrix}a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\\ 0&a_{1}&a_{3}\end{vmatrix}>0$, i.e., $\Delta_{k}>0$ for all $1\leq k\leq 3$. We need to show that the $4\times 4$ leading principal minor is positive, i.e.,
(1) $\frac{a_{1}}{a_{0}}>\frac{\begin{vmatrix}a_{3}&a_{5}&0\\ a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\end{vmatrix}}{\begin{vmatrix}a_{2}&a_{4}&0\\ a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\end{vmatrix}}=\frac{a_{3}-a_{5}\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}}{a_{2}-a_{4}\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}}.$
Let $r=\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}$. Since $\frac{a_{3}}{a_{2}}>\frac{a_{5}}{a_{4}}$, so by combining $(a)$ and $(c)$ of Proposition 4.8, we have $\frac{a_{3}-a_{5}\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}}{a_{2}-a_{4}\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}}>\frac{a_{3}}{a_{2}}$. By Lemma 4.17 the inequality in Equation 1 holds true if and only if
$\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}}$
$\Leftrightarrow$ $(a_{1}a_{4}-a_{0}a_{5})^{2}<(a_{1}a_{2}-a_{0}a_{3})(a_{3}a_{4}-a_{2}a_{5})$
In fact, a stronger result is presented in [KV08, Theorem 3] for entire functions. However, we limit our focus here to univariate polynomials, cf. [KV08, Theorem 1].
@@
If the coefficients of $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ are positive and satisfy the inequalities $a_{k}a_{k+1}\geq x_{0}a_{k-1}a_{k+2}$ for $11\leq k\leq d-$ 1 where $x_{0}$ is the unique positive root of the polynomial $x^{3}-x^{2}-2x-1(x_{0}\approx 2.1479)$, then $f(t)$ is Hurwitz-stable. In particular, the conclusion is true if $a_{k}^{2}\geq\sqrt{x_{0}}a_{k-1}a_{k+1}$ for $11\leq k\leq d-$ 1.
Theorem 4.20.
Let $f(x)\in{\mathbb{R}}[x]$ be a strictly ${\mathcal{K}}$-CLC of degree up to $4$ over a proper convex cone ${\mathcal{K}}$. Then $f(x)$ is a Hurwitz-stable polynomial over ${\mathcal{K}}$.
@@
Since $f$ is strictly ${\mathcal{K}}$-CLC, so by Definition 3.12, the univariate polynomial $f(x+tv)=f(x)+D_{v}f(x)t+\dots+\frac{D_{v}^{d}f(x)}{d!}t^{d}$ is a strictly CLC/Lorentzian for all $t$ such that $x+tv\in{\mathcal{K}}$. Moreover, for any $x,v\in{\mathcal{K}}$, all the coefficients of $f(x+tv)$ are positive by Corollary 3.14. Since $f(x)$ is strictly ${\mathcal{K}}$-CLC, so by Theorem 3.15, $D_{v}^{k}f(x)^{2}>D_{v}^{k-1}f(x)D_{v}^{k+1}f(x)$ for any $11\leq k\leq d-$ 1 and $x\in{\mathcal{K}}$. On the other hand, since the univariate restriction $f(x+tv)$ is Hurwitz-stable for any $x,v\in\operatorname{int}{\mathcal{K}}$, so by Definition 4.5, $f(x)$ is Hurwitz-stable polynomial over ${\mathcal{K}}$. Then the rest follows from Proposition 4.14. ∎
Theorem 4.21.
Let $f(x)\in{\mathbb{R}}[x]$ be a strictly ${\mathcal{K}}$-CLC of degree $d\geq 5$ over a proper convex cone ${\mathcal{K}}$. If $\left(D_{v}^{k}f(x)\right)^{2}>\frac{\alpha k}{k+1}D_{v}^{k-1}f(x)D_{v}^{k+1}f(x)$ for any $11\leq k\leq d-$ 1, $x,v\in\operatorname{int}{\mathcal{K}}$, and $\alpha$ is the unique positive root of the polynomial $t^{3}-t^{2}-2t-1(\alpha\approx 2.1479)$, then $f(x)$ is a Hurwitz-stable polynomial over ${\mathcal{K}}$.
Proof.
Since $f$ is strictly ${\mathcal{K}}$-CLC, so by Definition 3.12, the univariate polynomial $f(x+tv)=f(x)+D_{v}f(x)t+\dots+\frac{D_{v}^{d}f(x)}{d!}t^{d}$ is a strictly CLC/Lorentzian for all $t$ such that $x+tv\in{\mathcal{K}}$. Then the result follows from Theorem 4.19 by setting $f(x+tv)=g(t)=\sum_{k=0}^{d}a_{k}t^{d}$ for some $x,v\in{\mathcal{K}}$ and realizing $a_{k}=\frac{D_{v}^{k}(f)}{k!}$. ∎
@@
Not all ${\mathcal{K}}$-strictly CLC are Hurwitz stable polynomial over ${\mathcal{K}}$. For example, consider $f=5+14t+12.5t^{2}+7.2t^{3}+3t^{4}+t^{5}$. This polynomial is not Hurwitz-stable but Lorentzian for $t\geq 0$. On the other hand, consider $f=5+25t+50t^{2}+30t^{3}+10t^{4}+3t^{5}$. It is Hurwitz-stable but not Lorentzian.
5. Cone associated with ${\mathcal{K}}$-Lorentzian polynomials
We propose a method to construct a proper cone ${\mathcal{K}}$ for a given ${\mathcal{K}}$-Lorentzian polynomial. We define that $f$ is ${\mathcal{K}}$-Lorentzian with respect to $v\in\operatorname{int}{\mathcal{K}}$ if $f$ is Lorentzian along the ray passing through the point $v$. Note that the equivalence between the spaces of ${\mathcal{K}}$-Lorentzian polynomials and ${\mathcal{K}}$-CLC forms is established in [BD24, Theorem 4.10]).
In this context, we associate a proper cone ${\mathcal{K}}$ containing $v$ as an interior point of $f$, akin to the concept of hyperbolic polynomials with respect to $e$ and their corresponding hyperbolicity cone. We denote this cone as ${\mathcal{K}}(f,v)$, indicating its dependence on the polynomial $f$ and the point $v$.
First, we provide a polynomial inequality description of ${\mathcal{K}}(f,v)$, which is a semialgebraic set. Since $f$ is CLC with respect to $v\in\operatorname{int}{\mathcal{K}}$, all its directional derivatives along the direction of $v$ are log-concave. Therefore, by utilizing the Proposition 3.13 we define
$\operatorname{\mathring{{\mathcal{K}}}}(f,v)=\{x\in{\mathbb{R}}^{n}:f(x)>0,D_{v}f(x)>0,\dots,D_{v}^{d-1}f(x)>0\}$
By construction, $\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ is an open (semialgebraic) cone.
Proposition 5.1.
$\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ is a connected component of $\{x\in{\mathbb{R}}^{n}:f(x)\neq 0\}$ containing $v$.
Proof.
In order to show that $\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ is a connected component it’s enough to show there exists a path from $x$ to $v$ on which $f(x)$ remains strictly positive. Let $l$ be the line segment with endpoints $x,v$. For sufficiently large $\lambda$, all $y\in l$ satisfy $f(y+\lambda v)>0$. Since $x,v\in\operatorname{int}{\mathcal{K}}(f,v)$, due to linearity property of differentials, $x+tv,v+tv\in\operatorname{int}{\mathcal{K}}(f,v)$ for any $t\geq 0$. That implies that $f(x+tv)>0,f(v+tv)>0$ for any $t\geq 0$. Therefore, the segments $\{x+tv;0\leq t\leq\lambda\}$, $\{y+\lambda v;y\in l\}$ and $\{v+tv;0\leq t\leq\lambda\}$ form a path from $x$ to $v$ such that $f$ remains strictly positive. ∎
Remark 5.2.
Since for any $t>0$, $tv+w\in\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ for any $v,w\in\operatorname{\mathring{{\mathcal{K}}}}(f,v)$, so $\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ is a star-shaped set with respect to $v$. Since $v\in\operatorname{\mathring{{\mathcal{K}}}}(f,v)$, $\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ has a nonempty interior. Note that if $x\in\operatorname{\mathring{{\mathcal{K}}}}(f,v)$, then $-x\notin\operatorname{\mathring{{\mathcal{K}}}}(f,v)$. Since $\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ does not contain any line, so $\operatorname{\mathring{{\mathcal{K}}}}(f,v)$ is a proper star-shaped cone.
We define ${\mathcal{K}}(f,v)$ as follows
${\mathcal{K}}(f,v)=\{x\in{\mathbb{R}}^{n}:f(x)\geq 0,D_{v}f(x)\geq 0,\dots,D_{v}^{d-1}f(x)\geq 0\}$
Thus, ${\mathcal{K}}(f,v)$ is a closed set since it’s the finite intersection of closed semialgebraic sets. We show that ${\mathcal{K}}(f,v)$ is the Zariski closure of the connected component $\operatorname{\mathring{{\mathcal{K}}}}(f,v)$. The following results are required for the proof.
Proposition 5.3.
[BD24, Prop:4.5] Let $f\in{\mathbb{R}}[x]_{n}^{d}$ be a form of degree $d\geq 2$ and ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a proper convex cone. Suppose that $f(a)>0$ for all $a\in\operatorname{int}{\mathcal{K}}$, and let $Q_{a}$ denote the Hessian matrix of $f$ evaluated at a: $Q_{a}=H_{f}(a)$. The following statements are equivalent:

### v2

@@ 2. For $r_{1}$:
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ be a univariate strictly CLC polynomial for $t\geq 0$ with $a_{k}>0$ for $k=0,\dots,d$. If $1\leq d\leq 4$, then $f(t)$ is Hurwitz–stable.
@@
Assume $a_{0},a_{1},a_{2},a_{3},a_{4},a_{5}>0$ and $a_{2}-a_{4}k>0$. For any positive $k$,
$\frac{a_{1}}{a_{0}}>\frac{a_{3}-a_{5}k}{a_{2}-a_{4}k}\quad\Longleftrightarrow\quad k<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}},$
provided $a_{1}a_{4}-a_{0}a_{5}>0$.
@@
By the standard identity for positive denominators,
$\frac{a}{b}>\frac{c}{d}\;\Longleftrightarrow\;ad>bc.$
Applying this with $a=a_{1}$, $b=a_{0}$, $c=a_{3}-a_{5}k$, $d=a_{2}-a_{4}k$ and rearranging gives
$a_{1}(a_{2}-a_{4}k)>a_{0}(a_{3}-a_{5}k)\;\Longleftrightarrow\;(a_{1}a_{4}-a_{0}a_{5})k<a_{1}a_{2}-a_{0}a_{3},$
which is equivalent to the claimed inequality when $a_{1}a_{4}-a_{0}a_{5}>0$. ∎
Theorem 4.17.
Let $f(t)=\sum_{k=0}^{5}a_{k}t^{k}$ be a univariate polynomial with $a_{k}>0$ for $k=0,\dots,5$. Suppose that
$a_{k}^{2}>\frac{k+1}{k}\,a_{k-1}a_{k+1}\quad\text{for }k=1,\dots,4,$
and
$(a_{1}a_{4}-a_{0}a_{5})^{2}<(a_{1}a_{2}-a_{0}a_{3})(a_{3}a_{4}-a_{2}a_{5}).$
Then $f(t)$ is Hurwitz-stable.
@@
By Theorem˜4.2 and Remark˜4.3, for $d=5$ it is sufficient to show that all leading principal minors $\Delta_{k}>0,\quad k=1,\dots,4,$ of the Hurwitz matrix
${\mathrm{Hur}}_{f(t)}=\begin{bmatrix}a_{1}&a_{3}&a_{5}&0&0\\ a_{0}&a_{2}&a_{4}&0&0\\ 0&a_{1}&a_{3}&a_{5}&0\\ 0&a_{0}&a_{2}&a_{4}&0\\ 0&0&a_{1}&a_{3}&a_{5}\end{bmatrix}$.
The Newton-type inequalities $a_{k}^{2}>\frac{k+1}{k}\,a_{k-1}a_{k+1},\quad k=1,\dots,4,$ imply in particular that
$a_{1}>0,\quad a_{1}a_{2}-a_{0}a_{3}>0,\quad\det\begin{bmatrix}a_{1}&a_{3}&0\\ a_{0}&a_{2}&a_{4}\\ 0&a_{1}&a_{3}\end{bmatrix}>0.$
From this it follows that $\det\begin{bmatrix}a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\\ 0&a_{1}&a_{3}\end{bmatrix}>0$, so $\Delta_{k}>0$ for $k=1,2,3$.
It remains to show that the $4\times 4$ leading principal minor is positive, i.e.,
(2) $\frac{a_{1}}{a_{0}}>\frac{\det\begin{bmatrix}a_{3}&a_{5}&0\\ a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\end{bmatrix}}{\det\begin{bmatrix}a_{2}&a_{4}&0\\ a_{1}&a_{3}&a_{5}\\ a_{0}&a_{2}&a_{4}\end{bmatrix}}=\frac{a_{3}-a_{5}r}{a_{2}-a_{4}r},$
where $r:=\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}$. By the Newton-type inequalities and Proposition˜4.8(a),(c), we have
$\frac{a_{3}-a_{5}r}{a_{2}-a_{4}r}>\frac{a_{3}}{a_{2}}.$
On the other hand, by Lemma˜4.16, the inequality (2) holds if and only if
$r<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}},$
that is,
$\frac{a_{1}a_{4}-a_{0}a_{5}}{a_{3}a_{4}-a_{2}a_{5}}<\frac{a_{1}a_{2}-a_{0}a_{3}}{a_{1}a_{4}-a_{0}a_{5}}\;\Longleftrightarrow\;(a_{1}a_{4}-a_{0}a_{5})^{2}<(a_{1}a_{2}-a_{0}a_{3})(a_{3}a_{4}-a_{2}a_{5}),$
which is exactly our additional hypothesis. Hence the $4\times 4$ leading principal minor is positive and therefore $f(t)$ is Hurwitz-stable. ∎
In fact, a stronger result is presented in [Hurwitzsuff, Theorem 3] for entire functions. However, we limit our focus here to univariate polynomials, cf. [Hurwitzsuff, Theorem 1].
@@
If the coefficients of $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ are positive and satisfy
$a_{k}a_{k+1}\;\geq\;\alpha\,a_{k-1}a_{k+2},\qquad 1\leq k\leq d-1,$
where $\alpha$ is the unique positive root of the cubic
$t^{3}-t^{2}-2t-1=0\quad(\alpha\approx 2.1479),$
then $f(t)$ is Hurwitz-stable. In particular, the conclusion holds if
$a_{k}^{2}\;\geq\;\sqrt{\alpha}\,a_{k-1}a_{k+1},\qquad 1\leq k\leq d-1.$
Now we extend the above univariate results to multivariate homogeneous polynomials.
Theorem 4.19.
Let $f(x)\in{\mathbb{R}}^{d}_{n}[x]$ be a strictly $\mathcal{K}$}-Lorentzian polynomial of degree $d$ with $1\leq d\leq 4$ over a proper convex cone $\mathcal{K}$}. Then $f(x)$ is Hurwitz-stable over $\mathcal{K}$}.
@@
Since $f$ is strictly $\mathcal{K}$}-Lorentzian, Definition˜3.12 together with Theorem˜2.9 (the equivalence of strictly $\mathcal{K}$}-CLC and strictly $\mathcal{K}$}-Lorentzian polynomials in the homogeneous setting) implies that for any $x,v\in\operatorname{int}{\mathcal{K}}$ the univariate restriction
$f(x+tv)=\sum_{k=0}^{d}\frac{D_{v}^{k}f(x)}{k!}\,t^{k}$
is a strictly Lorentzian polynomial in $t$ for all $t$ such that $x+tv\in{\mathcal{K}}$. Moreover, by Corollary˜3.15, all coefficients $a_{k}:=\frac{D_{v}^{k}f(x)}{k!}$ are positive for $x,v\in\operatorname{int}{\mathcal{K}}$.
Since $f$ is strictly $\mathcal{K}$}-Lorentzian, Theorem˜3.16 yields the Newton-type inequalities
$\bigl(D_{v}^{k}f(x)\bigr)^{2}>\frac{k+1}{k}\,D_{v}^{k-1}f(x)\,D_{v}^{k+1}f(x),\qquad 1\leq k\leq d-1,$
for all $x,v\in\operatorname{int}{\mathcal{K}}$. In terms of the coefficients $a_{k}=\frac{D_{v}^{k}f(x)}{k!}$, this is exactly
$a_{k}^{2}>\frac{k+1}{k}\,a_{k-1}a_{k+1},\qquad 1\leq k\leq d-1.$
Therefore, by Proposition˜4.14, the univariate polynomial $t\mapsto f(x+tv)$ is Hurwitz-stable for every $x,v\in\operatorname{int}{\mathcal{K}}$. By Definition˜4.5, this means that $f(x)$ is Hurwitz-stable over $\mathcal{K}$}. ∎
@@
Let $f(x)\in{\mathbb{R}}^{d}_{n}[x]$ be strictly $\mathcal{K}$}-Lorentzian of degree $d\geq 5$ over a proper convex cone $\mathcal{K}$}. Suppose that
$\bigl(D_{v}^{k}f(x)\bigr)^{2}>\frac{\sqrt{\alpha}\,k}{k+1}\,D_{v}^{k-1}f(x)\,D_{v}^{k+1}f(x)\quad\text{for all }1\leq k\leq d-1,\;x,v\in\operatorname{int}{\mathcal{K}},$
where $\alpha$ is the unique positive root of the cubic $t^{3}-t^{2}-2t-1=0$ (so $\alpha\approx 2.1479$). Then $f(x)$ is Hurwitz-stable over $\mathcal{K}$}.
@@
Fix $x,v\in\operatorname{int}{\mathcal{K}}$ and consider the univariate restriction
$g(t):=f(x+tv)=\sum_{k=0}^{d}a_{k}t^{k},\qquad\text{with }a_{k}:=\frac{D_{v}^{k}f(x)}{k!}.$
By Definition˜3.12 and Theorem˜2.9, $g$ is strictly Lorentzian, so in particular $a_{k}>0$ for all $k$.
The assumed inequalities on $D_{v}^{k}f(x)$ translate into the coefficient inequalities required in Theorem˜4.18. Indeed,
$a_{k}^{2}=\frac{\bigl(D_{v}^{k}f(x)\bigr)^{2}}{(k!)^{2}}>\sqrt{\alpha}\,\frac{k}{k+1}\frac{D_{v}^{k-1}f(x)\,D_{v}^{k+1}f(x)}{(k!)^{2}}=\sqrt{\alpha}\,\frac{D_{v}^{k-1}f(x)}{(k-1)!}\,\frac{D_{v}^{k+1}f(x)}{(k+1)!}=\sqrt{\alpha}\,a_{k-1}a_{k+1}.$
Thus the sufficient condition in Theorem˜4.18 is satisfied, and so $g(t)$ is Hurwitz-stable. Since this holds for every $x,v\in\operatorname{int}{\mathcal{K}}$, it follows from Definition˜4.5 that $f(x)$ is Hurwitz-stable over $\mathcal{K}$}. ∎
@@
Not every strictly $\mathcal{K}$}-CLC polynomial is Hurwitz-stable over $\mathcal{K}$}. For instance, for ${\mathcal{K}}={\mathbb{R}}_{\geq 0}$, the univariate polynomial
$f(t)=5+14t+12.5t^{2}+7.2t^{3}+3t^{4}+t^{5}$
is Lorentzian for $t\geq 0$ but not Hurwitz-stable. Conversely, the polynomial
$f(t)=5+25t+50t^{2}+30t^{3}+10t^{4}+3t^{5}$
is Hurwitz-stable but not Lorentzian. These examples show that the sufficient conditions in Theorems˜4.19 and 4.20 are not necessary.
5. Applications: Stability Analysis for LTI Systems
Consider the linear time-invariant (LTI) system
$\dot{x}=Ax,\qquad A\in{\mathbb{R}}^{n\times n}.$
A matrix $A$ is called stable (or Hurwitz) if every eigenvalue of $A$ has negative real part, i.e., all eigenvalues lie in the open left half-plane. If $A$ is real symmetric, then $A$ is stable if and only if $A$ is negative definite. Lyapunov’s theorem asserts that $A$ is stable if and only if there exists a symmetric positive definite matrix $P$ such that the Lyapunov operator
$L_{A}(P):=A^{\top}P+PA$
is negative definite.
Theorem 5.1 (Routh–Hurwitz criterion).
Let $A\in{\mathbb{R}}^{n\times n}$, and let $\chi_{A}(t)$ be its characteristic polynomial. Then the following are equivalent [Poznyak, Chap. 9], [stabilityHurwitz, Chap. 7]: